How you you solve these logs finding x,b or y
logBASE3^y=2 and logBASEb^32=5?

Answer by stanbon(75887)

logBASE3^y=2 and logBASEb^32=5?
-------------
log3 y 2
y = 3^2 = 9
---------------
logb 32 = 5
b^5 = 32
b = 2
===============
Cheers,
Stan H.